OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1).

OpenStudy (anonymous):

parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (aroub):

parallel: y=mx+b , m=1 , (-2,-1) -1=1(-2)+b -1=-2+b 1=b =>y=x+1 perpendicular : y=mx+b , m=-1 , (-2,-1) -1=-1(-2)+b -1=2+b -3=b =>y=-x-3